x + 2y + 2z = 25
2x + y + z = 17
2x + 2y + z = 20

Answer by stanbon(75887)

x + 2y + 2z = 25
2x + y + z = 17
2x + 2y + z = 20
---------------------
Solve by any method you know to get:
x = 3
y = 3
z = 8
===========
Cheers,
Stan H.
===========